OpenStudy (anonymous):

give the next three numbers in the pattern 1,3,9,27,__,__

OpenStudy (anonymous):

sorry 3 numbers so ___,___,___,

OpenStudy (anonymous):

81, 243, 729

jhonyy9 (jhonyy9):

81,243,729

OpenStudy (anonymous):

You are basically multiply each term by 3,1*3=3, 3*9 = 27, 27*3= 81, 81*3=729